OpenStudy (anonymous):

Define relations R and S on B as follows: R={(x,y) ∈ B x B|x^2+y^2=4} and S={(x,y) ∈ B x B|x=y}

OpenStudy (anonymous):

Graph R, S the union of R and S and the intersection of R and S

OpenStudy (kinggeorge):

Are you given that B is? Or is it just some set?

OpenStudy (anonymous):

R is circle x^2+y+2=4 S is the line x=y  Union is both the figures and intersection is

OpenStudy (anonymous):

two points

OpenStudy (anonymous):

hahah well its not really B it really is the Bold R but I didnt want to confuse u guys

OpenStudy (kinggeorge):

In that case, follow @NotSObright's advice.

OpenStudy (anonymous):

Gotcha. Thanks that was great :)

OpenStudy (anonymous):

Yes, \(\mathbb R \times \mathbb R\) is the two-dimensional plane.